OpenStudy (anonymous):

write in the standard form of a hyperbola 9x^2 -3y^2 = 48y +192

zepdrix (zepdrix):

First step! divide through by the largest coefficient on either x or y. We need those out of the way :O What does your equation look like now? :)

OpenStudy (anonymous):

I'm lost ;/

OpenStudy (anonymous):

\[\frac{ 3x ^{2} }{ 16 } - \frac{ \left( y+4 \right) ^{2}}{ 48 } = 1\]